OpenStudy (anonymous):

If a car travels 180 miles in 3 hours, how far would it travel in 5 hours? Select one: a. 200 b. 225 c. 275 d. 300

OpenStudy (anonymous):

@Sofitina

OpenStudy (anonymous):

Math is not my forte but I think it's 300.

OpenStudy (anonymous):

She's right it's 300 to do this problem use the formula\[\frac{ distance }{ time }\] then multiply the outcome by 5 \[\frac{ 180 }{ 3 }= 60×5=300\]

OpenStudy (anonymous):

I'm proud of myself B)